OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (anonymous):

solve for y and the x the substitute

OpenStudy (michele_laino):

please add to both sides of the second equation -7, then write what you got

OpenStudy (michele_laino):

for example if I add -7 to both sides of the second equation, then the left side of that equation becomes: y-7, and the right side? please try

OpenStudy (anonymous):

y = -2x + 19 y = x + 7 answer choices below : x = 4, y = 11 x = -12, y = -5 x = -4, y = 3 x = 8, y = 3

OpenStudy (michele_laino):

sorry I can not say the solution directly since the code of conduct, nevertheless I can help you in your reasoning